on a certain day ,rain was falling vertcally with a speed of 30m/s .A wind started blowing after some time with a speed of 10m/s in north-south dirction . In which direction dhould a boy waiting at a bus stop hold his umbrella to protect himself from rain?

Grade:11

1 Answers

CHEETALI SHARMA
2 years ago
The velocity of wind and velocity of rain are in same order and resultant will be in opposite order
wind =30 ms-1
rain = 10 ms-1
Resultant velocity is somewhere between the V wind and V rain at some angle theta
Tan theta = V rain / V wind
Tan theta=10/30= 1/3
Theta = tan-1 1/3